A man invests $2,200 in three accounts that pay 6%, 8%, 9% in annual interest, respectively. he has three times as much invented at 9% as he does at 6%. if his total interest for year is $178, how much is invested at each rate?
Answer by mananth(16946)

6% --- amount x
9%---- amount 3x
8% --- amount 2200- (3x+x)
..
Interest recieved = 178
..
0.06x+0.09*3x+0.08(2200-4x)=178
0.06x+0.27x+176.00-0.32x=178
0.01x=2
x=200 at6%
3x= 600 at9%
2200-4x= 2200-800=1400 at 8%